Fermi National Laboratory


Accelerator Update

Friday June 18
The day shift began with Operations monitoring store 3582 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Operations resumed sending beam to the Meson MTest experiment at 10 AM.

Recycler began stashing antiprotons at 11:04 AM.

Pbar resumed stacking at 12:20 PM.

Operations established beam to the Meson MCenter at 7:45 PM.

Saturday June 19
The midnight shift began with Operations monitoring store 3582 and the stash, with Pbar stacking, and with MiniBooNE and MCenter taking beam.

At 12:34 AM, a Booster RF station (BRF12) tripped off for the first of two times in five minutes.

At 1:43 AM, a Linac RF station (KRF7) tripped off and would not reset. Operators investigated and then called in a Linac technician.

At 4:03 AM, another Booster RF station (BRF18) tripped off. It reset okay.

A Linac tech reported at 4:30 AM that KRF7 was fixed. Pbar resumed stacking.

At 5:48 AM, a Linac RF station (KRF6) tripped off on a water flow indication. It reset okay, but it took five minutes for the station to resume its normal phase.

At 6:41 AM, Operations reported that a Linac RF station had tripped off twice. The station reset itself both times.

At 8:24 AM, a Linac RF station (KRF2) tripped off. Operators investigated and discovered a power supply problem. Operators first threatened the power supply and when that didn't work, smacked it a couple of times. The supply straightened up and began running normally.

Tevatron (TeV) experts began an end-of-store study at 10:30 AM. At 11:35 AM, TeV experts began an abort-gap study.

Operations terminated store 3582 at 12:20 PM.

Operations put the TeV through a dry squeeze at 12:54 PM.

BRF18 tripped off at 1:05 PM. It reset okay.

Operations put the TeV into shot setup at 1:22 PM.

Operations established store 3584 at 3:44 PM with an initial luminosity of 77.66E30. Pbar resumed stacking.

At 4:11 PM, Operations reported the TeV had a colliding beam problem. An expert suggested a full collimator move. Operators actually did the procedure twice and then put Pbar back into stacking mode at 4:22 PM.

LRF4 tripped off at 4:53 PM. It reset itself okay.

BRF18 tripped off at 6:58 PM on a power supply overload. It reset okay.

Operators made Recycler ready to stash antiprotons at 7:28 PM. They completed their stash and Pbar resumed stacking 8:34 PM.

Pbar experts began a zero-stack study at 9 PM.

Sunday June 20
The midnight shift began with Operations monitoring store 3584 and the stash, with Pbar experts conducting studies, and with MiniBooNE, MTest, and MCenter taking beam.

Operations reported at 12:02 AM that the QXR wasn't operating normally. Operators tried various reboots with no success. Operators contacted an expert who determined that a slow spill bit had been switched to the off position during a reboot.

At 6:56 AM, a Booster power supply (EAPS) had tripped off due to problems with BRF2. Operator investigated and then called in a tech to look at BRF2. Operators bypassed the station at 9 AM so Pbar could resume stacking.

BRF12 tripped off at 9:54 AM. It reset okay.

The Booster RF tech reported that he replaced a modulator for BRF2. Operators had it back on line by 10:49 AM. Normal beam resumed at 11:11 AM.

BRF12 tripped off at 11:34 AM. It reset okay.

MTest experiments said at 1:37 PM that they were finished with beam.

BRF12 tripped off at 3:20 PM. It reset okay.

TeV experts began an end-of-store study at 8 PM.

At 9:12 PM, the Main Injector (MI) loop tripped off on a possible ground fault. Operators conducted a hipot and found no problems. Then both MI and MiniBooNE tripped off at 9:16 PM due to a beam problem. Operators discovered two trim magnets power supplies had tripped off. Operator reset the supplies and normal beam resumed.

BRF12 tripped off at 9:41 PM. It reset okay.

TeV expert completed the end-of-store study and Operations terminated store 3584 at 9:57 PM.

Operation put the TeV through a dry squeeze and then into shot setup at 10:48 PM.

Monday June 21
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, with MiniBooNE taking beam, and with MCenter off due to shot setup.

Operations established store 3586 at 12:56 AM with an initial luminosity of 82.8E30. This set a NEW LUMINOSITY RECORD.

Pbar resumed stacking until a Recycler expert arrived at 3:31 AM to begin stashing antiprotons.

Recycler completed stashing antiprotons at 4:56 AM. Operators resumed sending beam to MCenter.

Operations reported finding a Pbar Debuncher trim magnet power supply off at 6:40 AM. The stacking rate looked okay. Operators investigated and found a bad fuse. They replaced the fuse, turned on the supply, and the stacking rate increased a little.

Plans for Today
The plans for today are to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/21/2004   email Fermilab